A two point system consists of two masses, m and 2m, separated by a distance d. If mass m is replaced by a mass 3m and is kept at a distance d/2, what are the ratios of the gravitational potential energy of the two systems?


1
Expert's answer
2021-11-24T11:58:43-0500

"F_1=\frac{k\cdot m\cdot 2m}{d^2}=\frac{2km^2}{d^2},"

"F_2=\frac{k\cdot 3m\cdot 2m}{(\frac d2)^2}=\frac{24m^2}{d^2},"

"\frac{F_2}{F_1}=12."